• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name folts

    Log entry time 16:47:16 on June 01, 2010

    Entry number 324115

    keyword=target vacuum/tech

    The target chamber has been re-connected to the spectrometers and is pumping down.

    The transition O rings are silicone just like they were before and the beam line O ring is the 7075 kelraz

    If all goes well we will do a leak check and alignment check in the morning and if the vacuum is good enough possibly cool down in the late afternoon